# Interior structure of black holes with nonlinear terms

### v1

@@ I Introduction
As one of the most significant astrophysical objects in gravitational physics, black holes remain at the forefront of fundamental research. On the one hand, treating black holes as thermodynamic systems allows us to explore profound connections between quantum gravity and statistical physics Bekenstein:1973ur; Hawking:1975vcx. On the other hand, based on the AdS/CFT correspondence Maldacena:1997re, the holographic properties of black holes provide a powerful and unique theoretical framework for investigating phase transitions such as high-temperature superconductivity Hartnoll:2008vx; Hartnoll:2008kx; Herzog:2010vz; Gubser:2008wv; Cai:2013aca; Chen:2010mk; Kim:2013oba; Basu:2010fa; Musso:2013ija; Nie:2013sda; Amado:2013lia; Donos:2013woa; Nie:2015zia; Nie:2014qma; Li:2017wbi; Zhang:2021vwp, strongly coupled quantum field theories and particularly non-equilibrium processes in condensed matter physics Zeng:2022hut; Xia:2023pom; Xia:2021jzh; Zhao:2023ffs; Su:2023vqa; Zeng:2024rwn.
@@
For a long time, research in black hole physics has predominantly focused on their astrophysical properties. However, a more fundamental and fascinating question arises: what physical laws govern the internal structure of black holes beyond the event horizon? Recent studies have shown that scalarized black holes exhibit rich dynamical behaviors in their interiors Cai:2020wrp; Hartnoll:2020fhc; Wang:2020nkd; An:2021plu; Mansoori:2021wxf; Liu:2021hap; Cai:2021obq; An:2022lvo; Sword:2022oyg; Xu:2023fad; Zhang:2025hkb; Zhang:2025tsa; Xu:2025edz: the inner horizon disappears and is replaced by a series of dynamical regimes, including the collapse of the Einstein-Rosen bridge, Josephson oscillations, ultimately settling into a Kasner universe. The behavior of the scalar field in the Kasner regime is characterized by the Kasner exponent $p_{t}$, which exhibits highly oscillatory behavior near the critical point of the scalarization phase transition.
@@
A natural question follows: does this oscillatory structure, which encodes information about the black hole interior, possess deeper regularities? Can its pattern be precisely described and controlled? Meanwhile, introducing nonlinear interaction terms into holographic superfluid models has been shown to significantly alter the system’s phase transition and condensation behavior Zhao:2022jvs; Zhao:2024jhs; Zhao:2025tqq. This motivates us to investigate: could nonlinear terms also exert a crucial influence on the internal structure of black holes, particularly on the oscillatory behavior of the Kasner exponent?
@@
This work aims to systematically study the impact of higher-order nonlinear terms on the interior Kasner geometry of black holes within a holographic s-wave superfluid model. Our numerical calculations find that the nonlinear coefficient $\lambda$ can effectively tune the period of the oscillatory behavior: a positive $\lambda$ stretches the region, while a negative $\lambda$ compresses it. This discovery not only provides a new perspective for understanding the complex periodic structure within black holes but also reveals a novel approach to actively controlling this structure through model parameters.
@@ II Holographic setup
We first briefly introduce the model we employ. We consider a charged scalar field with additional higher-order nonlinear terms $\lambda(\psi^{\ast}\psi)^{2}$ and $\tau(\psi^{\ast}\psi)^{3}$. When the coefficients of nonlinear term $\lambda$ and $\tau$ are equal to zero, it reduces to the simplest holographic s-wave superconductor model Hartnoll:2008vx; Hartnoll:2008kx. Thus, the total action takes the following form
@@ III Interior solutions of black holes and Kasner geometry
The addition of an extra charged scalar field to the black hole system leads to the disappearance of the inner Cauchy horizon and the formation of a series of dynamical regions Cai:2020wrp; Hartnoll:2020fhc. These regions include the collapse of the ER bridge, followed by the Josephson oscillation, and finally the Kasner universe. In some cases, Kasner inversion, transition, alternation behavior may also occur. Detailed analytical analysis has been provided in Ref Hartnoll:2020fhc.
Before diving into numerical analysis, we first need to obtain the explicit form of the Kasner metric. Following Ref. Hartnoll:2020fhc, we can neglect the mass term of the scalar field and the charge term in the Maxwell equation. By ignore the coefficients of the higher-order nonlinear terms of scalar field. The Eqs. (8) (11) then become
@@
Using Eq. (22), the solutions of Eqs. (21) (24) can be written as
$f$ $=-f_{K}z^{3+\alpha^{2}}+\cdots~,$
$\psi$ $=\alpha\sqrt{2}\log z+\cdots~,$
$\chi$ $=2\alpha^{2}\log z+\chi_{K}+\cdots~,$ (25)
@@
Transforming the z coordinate to the proper time $\tau$, we can obtain the Kasner form of the metric
$ds^{2}$ $=-d\tau^{2}+c_{t}\tau^{2p_{t}}dt^{2}+c_{x}\tau^{2p_{x}}(dx^{2}+dy^{2})~,$
$\psi$ $=-p_{\psi}\log\tau~.$ (27)
where
@@ IV The internal periodical structure of black hole
As studied in Refs. Zhang:2021vwp; Zhao:2022jvs; Zhao:2023ffs; Zhao:2024jhs; Cao:2024irr; Zhao:2025tqq, when the extra higher-order nonlinear terms are added to the system, both the condensation behavior and the free energy of the system will change accordingly. However, the aforementioned scenario only considers the influence of the nonlinear terms on the condensation value and does not involve the internal structure of the black hole. In this paper, we will focus specifically on the effect of the nonlinear terms on the internal structure of the black hole.
First, we briefly review the influence of the nonlinear term on the system. In Fig. 1, we present the condensation curves for different values of the parameter $\lambda$ with $\tau=0$. As previously studied in Ref. Zhao:2025tqq, due to the back-reaction effect, even with a negative $\lambda$, the system does not exhibit a zeroth-order phase transition. In Fig. 1, the influence of the nonlinear term on the condensation curve is limited to modifying the value of the condensate. Meanwhile, in Fig. 2, we show the interior solution of the black hole for different values of $\lambda$ in $T=0.981T/T_{c}$. By computing the Kasner exponents for all temperatures, we obtain the relationship between $p_{t}$ and temperature $T$, as illustrated in Fig. 3.
@@
Figure 2: The Josephson oscillations behavior for $T=0.981T/T_{c}$ with $\lambda=$0.4, 0 and 0.4-0.4.
@@
As mentioned in Ref. Hartnoll:2020fhc, the Kasner exponents $p_{t}$ exhibit highly oscillatory behavior near the critical point. In this paper, by adding a higher-order nonlinear term, we can stretch or compress the interval of this oscillatory behavior. For example, in Fig. 3, when $\lambda$=0 (cyan curve), the oscillatory interval is very close to the critical point, and as the temperature decreases, the system will encounter a second oscillatory interval. However, when $\lambda$=0.4 (red curve), the oscillatory interval is significantly enlarged, and the second oscillatory interval will disappear. In contrast, when $\lambda=-0.4$ (green curve), the oscillatory interval is further compressed near the critical point. Therefore, $\lambda$ can be regarded as a magnifying or reducing factor to control the size of the oscillatory region of the Kasner exponents $p_{t}$.
In addition, this periodic behavior is a function of the form $sin(1/x)$, and its structure can be represented by an analytically solvable expressionHartnoll:2020fhc
@@ V Conclusions
Acknowledgements
We are grateful to Xin Zhao for useful discussions. This work is supported by National Natural Science Foundation of China (Grant Nos. 12533001, 12575049, 12473001, 12205039, 12305058, 11965013, 12575054 and 12405066). ZYN is partially supported by Yunnan High-level Talent Training Support Plan Young &\& Elite Talents Project (Grant No. YNWR-QNBJ-2018-181). This work is also supported by the National SKA Program of China (grant Nos. 2022SKA0110200 and 2022SKA0110203) and the 111 Project (Grant No. B16009).

### v2

@@ I Introduction
As one of the most significant astrophysical objects in gravitational physics, black holes remain at the forefront of fundamental research. On the one hand, treating black holes as thermodynamic systems allows us to explore profound connections between quantum gravity and statistical physics Bekenstein (1973); Hawking (1975). On the other hand, based on the AdS/CFT correspondence Maldacena (1998), the holographic properties of black holes provide a powerful and unique theoretical framework for investigating phase transitions such as high-temperature superconductivity Hartnoll et al. (2008a, b); Herzog (2010); Gubser and Pufu (2008); Cai et al. (2014); Chen et al. (2010); Kim and Taylor (2013); Basu et al. (2010); Musso (2013); Nie et al. (2013); Amado et al. (2014); Donos et al. (2014); Nie and Zeng (2015); Nie et al. (2015); Li et al. (2018); Zhang et al. (2022), strongly coupled quantum field theories and particularly non-equilibrium processes in condensed matter physics Zeng et al. (2023); Xia et al. (2023, 2022); Zhao et al. (2024); Su et al. (2024); Zeng et al. (2025).
@@
For a long time, research in black hole physics has predominantly focused on their astrophysical properties. However, a more fundamental and fascinating question arises: what physical laws govern the internal structure of black holes beyond the event horizon? Recent studies have shown that scalarized black holes exhibit rich dynamical behaviors in their interiors Cai et al. (2021); Hartnoll et al. (2021); Wang et al. (2020); An et al. (2021); Mansoori et al. (2021); Liu et al. (2021); Cai et al. (2022); An et al. (2022); Sword and Vegh (2022); Xu et al. (2024); Zhang et al. (2025a, b); Xu et al. (2025): the inner horizon disappears and is replaced by a series of dynamical regimes, including the collapse of the Einstein-Rosen bridge, Josephson oscillations, ultimately settling into a Kasner universe. The behavior of the scalar field in the Kasner regime is characterized by the Kasner exponent $p_{t}$, which exhibits highly oscillatory behavior near the critical point of the scalarization phase transition.
@@
A natural question follows: does this oscillatory structure, which encodes information about the black hole interior, possess deeper regularities? Can its pattern be precisely described and controlled? Meanwhile, introducing nonlinear interaction terms into holographic superfluid models has been shown to significantly alter the system’s phase transition and condensation behavior Zhao et al. (2023, 2025a, 2025b). This motivates us to investigate: could nonlinear terms also exert a crucial influence on the internal structure of black holes, particularly on the oscillatory behavior of the Kasner exponent?
@@
This work aims to systematically study the impact of higher-order nonlinear terms on the interior Kasner geometry of black holes within a holographic s-wave superfluid model. Our numerical calculations find that the nonlinear coefficient $\lambda$ can effectively tune the period of the oscillatory behavior: a positive $\lambda$ stretches the region, while a negative $\lambda$ compresses it. This discovery not only provides a new perspective for understanding the complex periodic structure within black holes but also reveals a novel approach to actively controlling this structure through model parameters. The behavior inside a black hole is indeed rich, and a complete description involves multiple regions extending from the horizon to the singularity. The Kasner exponents provide a relatively simple quantitative measure and exhibit interesting inverse periodic behavior near the critical point. Typically, universal laws near a critical point hold only within an infinitesimally small linear region around it. However, this inverse periodic behavior remains stable over a finite range near the critical point, making it particularly intriguing to explore the influence of nonlinear terms on this phenomenon. Such investigation can deepen our understanding of the critical universality exhibited inside black holes.
@@ II Holographic setup
We first briefly introduce the model we employ. We consider a charged scalar field with additional higher-order nonlinear terms $\lambda(\psi^{\ast}\psi)^{2}$ and $\tau(\psi^{\ast}\psi)^{3}$. When the coefficients of nonlinear term $\lambda$ and $\tau$ are equal to zero, it reduces to the simplest holographic s-wave superconductor model Hartnoll et al. (2008a, b). Thus, the total action takes the following form
@@ III Interior solutions of black holes and Kasner geometry
The addition of an extra charged scalar field to the black hole system leads to the disappearance of the inner Cauchy horizon and the formation of a series of dynamical regions Cai et al. (2021); Hartnoll et al. (2021). Beyond the horizon, the metric component $g_{tt}$ undergoes a rapid exponential contraction, a process that destroys the original inner horizon (the collapse of ER bridge). Subsequently, driven by the background gauge potential $A_{t}$, the charged scalar field exhibits behavior analogous to the Josephson effect in superconductivity. These oscillations leave observable imprints on the metric functions and serve as a key transitional phase connecting the initial collapse to the later Kasner universe; their specific parameters determine the subsequent evolution of spacetime (Josephson oscillation). After the Josephson oscillation, the spacetime geometry settles into a stable Kasner epoch. In this Kasner regime, all spatial metric components evolve as power laws in time (here, either the $z$-coordinate inward from the horizon or the proper time $\xi$), while the scalar field evolves logarithmically. Moreover, the Kasner exponents satisfy the relations that their sum and the sum of their squares both equal one. Ultimately, the final Kasner epoch terminates at the singularity. In some cases, Kasner inversion, transition, alternation behavior may also occur. Detailed analytical analysis has been provided in Ref Hartnoll et al. (2021).
Before diving into numerical analysis, we first need to obtain the explicit form of the Kasner metric. Following Ref. Hartnoll et al. (2021), we can neglect the mass term of the scalar field and the charge term in the Maxwell equation. By ignoring the the higher-order nonlinear terms of scalar field. The Eqs. (8) (11) then become
@@
Using Eq. (22) and setting the scalar field as $\psi=\sqrt{2}\int(\Psi/z)dz$, where $\Psi=\alpha$ (analytic solution to Eq. (22) for the scalar field; see Ref. Hartnoll et al. (2021) for details), the solutions of Eqs. (21) (24) can be written as
$f$ $=-f_{K}z^{3+\alpha^{2}}+\dots~,$
$\psi$ $=\alpha\sqrt{2}\log z+\dots~,$
$\chi$ $=2\alpha^{2}\log z+\chi_{K}+\dots~,$ (25)
@@
Transforming the z coordinate to the proper time $\xi$, we can obtain the Kasner form of the metric
$ds^{2}$ $=-d\xi^{2}+c_{t}\xi^{2p_{t}}dt^{2}+c_{x}\xi^{2p_{x}}(dx^{2}+dy^{2})~,$
$\psi$ $=-p_{\psi}\log\xi~.$ (27)
Here $c_{t}$ and $c_{x}$ are constants and
@@ IV The internal periodical structure of black hole
As studied in Refs. Zhang et al. (2022); Zhao et al. (2023, 2024, 2025a); Cao et al. (2025); Zhao et al. (2025b), when the extra higher-order nonlinear terms are added to the system, both the condensation behavior and the free energy of the system will change accordingly. However, the aforementioned scenario only considers the influence of the nonlinear terms on the condensation value and does not involve the internal structure of the black hole. In this paper, we will focus specifically on the effect of the nonlinear terms on the internal structure of the black hole.
First, we briefly review the influence of the nonlinear term on the system. In Fig. 1, we present the condensation curves for different values of the parameter $\lambda$ with $\tau=0$. As previously studied in Ref. Zhao et al. (2025b), due to the back-reaction effect, even with a negative $\lambda$, the system does not exhibit a zeroth-order phase transition. In Fig. 1, the influence of the nonlinear term on the condensation curve is limited to modifying the value of the condensate. Meanwhile, in Fig. 2, we show the interior solution of the black hole for different values of $\lambda$ in $T=0.981T/T_{c}$. By computing the Kasner exponents for all temperatures, we obtain the relationship between $p_{t}$ and temperature $T$, as illustrated in Fig. 3.
@@
Figure 2: The interior behavior of scalar filed for $T=0.981T/T_{c}$ with $\lambda=$0.4, 0 and 0.4-0.4. The upper panel illustrates the complete behavior of the scalar field inside the black hole, while the lower panel corresponds to the collapse of the ER bridge.
@@
As mentioned in Ref. Hartnoll et al. (2021), the Kasner exponents $p_{t}$ exhibit highly oscillatory behavior near the critical point. In this paper, by adding a higher-order nonlinear term, we can stretch or compress the interval of this oscillatory behavior. For example, in Fig. 3, when $\lambda$=0 (cyan curve), the oscillatory interval is very close to the critical point, and as the temperature decreases, the system will encounter a second oscillatory interval. However, when $\lambda$=0.4 (red curve), the oscillatory interval is significantly enlarged, and the second oscillatory interval will disappear. In contrast, when $\lambda=-0.4$ (green curve), the oscillatory interval is further compressed near the critical point. Therefore, $\lambda$ can be regarded as a magnifying or reducing factor to control the size of the oscillatory region of the Kasner exponents $p_{t}$.
In addition, this periodic behavior is a function of the form $sin(1/x)$, and its structure can be represented by an analytically solvable expressionHartnoll et al. (2021)
@@ V Conclusions
Acknowledgements We are grateful to Xin Zhao for useful discussions. This work is supported by National Natural Science Foundation of China (Grant Nos. 12533001, 12575049, 12473001, 12205039, 12305058, 11965013, 12575054 and 12405066). ZYN is partially supported by Yunnan High-level Talent Training Support Plan Young &\& Elite Talents Project (Grant No. YNWR-QNBJ-2018-181). This work is also supported by the National SKA Program of China (grant Nos. 2022SKA0110200 and 2022SKA0110203) and the 111 Project (Grant No. B16009).
Data Availability Statement This manuscript has associated data in a data repository. [Authors’ comment: The datasets generated during the current study are available in Zenodo repository, https://doi.org/10.5281/zenodo.18238745]
Code Availability Statement This manuscript has no associated code/software. [Author’s comment: Code/Software sharing not applicable to this article as no code/software was generated or analysed during the current study.]